OpenStudy (anonymous):

i need this answer quick. which is a solution of the equation? y = 4x + 3? a. 5,17 b. 17,5 c. 4,19 d. 4,16

OpenStudy (haleyelizabeth2017):

C

OpenStudy (haleyelizabeth2017):

you just need to plug in all of the points given and only one answer will be true

OpenStudy (anonymous):

can you help with another @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Of course! :D

OpenStudy (anonymous):

solve the equation. 2x - 6 = 2(x - 8) + 10 a. x = 2 b. x = 8 c. no solution d. infinitely many solutions

OpenStudy (haleyelizabeth2017):

okay one second.

OpenStudy (haleyelizabeth2017):

2x-6=2x-16+10 2x-6=2x-6 It will be d because whatever value you plug in for x, it will always be the same :)

OpenStudy (haleyelizabeth2017):

I think.....I could be wrong lol